What is the magnetic moment associated with a coil of one turn,are of a cross section 10^-4 m^2 carrying current of 2A

Grade:12

1 Answers

Khimraj
7 years ago
Magnetic moment M = NIA . So M =1*2*10^-4 = 2*10^-4 amp m².Hope it clears. If still u have doubts then please clarify. If u like answer then please approve the answer.